[Brahms-l] Shift report 20040119 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Mon Jan 19 2004 - 00:02:57 EST
------------------------------------------------------------
Date: Mon Jan 19 00:02:56 2004, Shift: 16:00-24:00
Supervisor: CEJ, with: 
Log book#15 pages 36 - 36

Shift Summary:
------------------------------------------------------------
1) Beam on untill 19:30. Again from 21:15.

2) Pedestal runs: 9752, 9753
   Zero field runs: 9749, 9750
   Data runs: FS@12 D1=A861, MRS@90, D5=B2255
   9745 : nTr6=100K , nTr3=55K
   9746 :      100K ,      32K
   9747 :      100K ,      31K
   9748 :       83K ,      26K
   Data runs: FS@12 D1=A861, MRS@90, D5=B2255
   9751 : nTr6= 80K , nTr3=21K
   9754 :       36K ,      23K
   9755 :       57K ,      26K
   9757 :      100K ,      26K
   9758 :      100K ,      26K
   9759 : still running.

3) The usual DC problems. T3 A2-xu was off during shift.
   T1 and T2 have a few noisy pads (lines in tpc3d).
   The meas current on the magnets are off by 0.5-2%.
   D2 meas current is 16 amps, but it's ok (magnet support).
   H1 slat 16 top TDC is dead.

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Jan 19 00:03:08 2004

This archive was generated by hypermail 2.1.8 : Mon Jan 19 2004 - 00:03:27 EST